### v1

# references.bib

Abstract
Kernel-free quadratic surface support vector machine (SVM) models have gained significant attention in machine learning. However, introducing a quadratic classifier increases the model’s complexity by quadratically expanding the number of parameters relative to the dimensionality of the data, exacerbating overfitting. To address this, we propose sparse $\ell_{0}$-norm based Kernel-free quadratic surface SVMs, designed to mitigate overfitting and enhance interpretability. Given the intractable nature of these models, we present a penalty decomposition algorithm to efficiently obtain first-order optimality points. Our analysis shows that the subproblems in this framework either admit closed-form solutions or can leverage duality theory to improve computational efficiency. Through empirical evaluations on real-world datasets, we demonstrate the efficacy and robustness of our approach, showcasing its potential to advance Kernel-free quadratic surface SVMs in practical applications while addressing overfitting concerns. All the implemented models and experiment codes are available at https://github.com/raminzandvakili/L0-QSVM.
keywords: binary classification, quadratic support vector machines, $\ell_{0}$ norm, penalty decomposition method
1 Introduction
Binary classification, a cornerstone of supervised learning, has garnered significant interest across various disciplines, including computer science, statistics, engineering, biology, and applied mathematics [di2007survey, langley1995applications]. Its practical applications in fields such as healthcare, finance, retail, travel, and energy highlight its widespread importance [hao2007new, monostori1996machine, luo2018benchmarking, chen2021equivalence]. Despite its evident utility, there remains a pressing need for rigorous theoretical scrutiny and validation of existing methodologies. To address the challenges of binary classification, the soft margin SVM model [cortes1995support] has emerged as a powerful tool, capable of finding hyperplanes that maximize margins while minimizing misclassification penalties. However, its effectiveness diminishes when dealing with nonlinearly separable data. To overcome this limitation, kernel based SVM models were introduced [cortes1995support]. These models leverage nonlinear kernel functions to map data into higher-dimensional feature spaces, enabling improved performance. Nevertheless, selecting an appropriate kernel function remains a challenge due to the lack of a general principle, and tuning kernel parameters involves significant computational overhead during training [cristianini2000introduction, scholkopf2001learning].
To maintain the interpretability of SVMs while sidestepping the complexities of nonlinear kernel methods, kernel-free quadratic surface SVM models have emerged. These models aim to find quadratic separating surfaces that maximize approximations of geometric margins while penalizing misclassifications of outliers and noise. As a result, they are more tractable than kernel based SVM models due to the explicit and clear quadratic structure of the separation surface [luo2016soft]. Additionally, they can incorporate prior information, such as Universum points, to further enhance performance. Recent advancements in this field have introduced twin versions of these models, addressing computational challenges by solving decoupled subproblems more efficiently [moosaei2023sparse, moosaei2024kernel].
Despite their advantages, kernel-free quadratic surface SVM models, when applied to $n$-dimensional data points, require $\mathcal{O}(n^{2})$ parameters, introducing risks of overfitting, redundant computations, and reduced interpretability, even when $n$ is moderately large. $\ell_{1}$-regularization of the Hessian has been employed to mitigate these issues by promoting sparsity in the parameter space. However, selecting the optimal penalty parameter remains a challenging task, and the solutions obtained may lack clear interpretability. To address these limitations, we propose $\ell_{0}$-norm based kernel-free quadratic surface SVMs that explicitly enforce sparsity in the model parameters. This approach enhances the explainability of the model by implicitly selecting critical features from the data. By combining the strengths of kernel-free quadratic surface methods with sparsity constraints, our technique offers a promising solution to the challenges associated with traditional kernel-free quadratic surface SVM models, paving the way for more robust and interpretable binary classification frameworks.
Penalty decomposition algorithms are well-regarded for their ability to address nonconvex optimization problems by decomposing them into simpler subproblems that are computationally manageable [lin2025inexact, lu2022penalty, mousavi2023cardinality, mousavi2024statistical]. Building on these strengths, we developed an effective penalty decomposition algorithm tailored to our proposed sparse $\ell_{0}$-norm based kernel-free quadratic surface SVMs. This approach not only addresses the inherent intractability of such models but also ensures the enforcement of sparsity constraints, enabling improved model interpretability and practical applicability. We further demonstrated that the subproblems within our algorithm are either solved in closed form or through duality theory to improve computational efficiency.
The rest of the paper is organized as follows. Section 2 provides an overview of related work, including existing methods and their limitations. Section 3 introduces the proposed sparse $\ell_{0}$-norm based kernel-free quadratic surface SVM models and the penalty decomposition algorithm. In Section 4, we present numerical experiments on real-world datasets to evaluate the performance of the proposed models, highlighting key insights and comparisons with baseline methods. Finally, Section 5 concludes the paper with a summary of contributions and potential directions for future work.
2 Preliminaries and Related Work
This section reviews several quadratic surface support vector machine (QSVM) models shortly. Suppose we have a binary classification task in hand with the following training set:
$\mathcal{D}=\{\left({x_{i}},y_{i}\right)_{i=1,\dots,m}\ \left|\ {x_{i}}\,\in\,\mathbb{R}^{n},\ y_{i}\in\{-1,1\}\right.\},$
where $m$ is the sample size, $n$ is the number of features, and $y_{i}$ is the label of $x_{i}$. $\mathcal{D}$ is called quadratically separable [luo2016soft] if there exist $W\in\mathbb{R}^{n\times n},b\in\mathbb{R}^{n},\,\text{and}\,c\in\mathbb{R}$ if $y_{i}(x_{i}^{T}W{x_{i}}+b^{T}{x_{i}}+c)>0$ for every $i\in[m]$. A linearly separable data set is quadratically separable with $W=0$.
The standard SVM model aims to find a hyperplane that approximately separates a given training set. However, in real-world binary classification problems, datasets often exhibit nonlinear patterns, making separation with traditional linear approaches difficult. To overcome this, kernel methods are employed to map the original data into a higher-dimensional feature space, where nonlinear relationships can appear linear. This transformation enables SVMs to find optimal hyperplanes that separate the data more effectively, capturing complex structures and patterns. As a result, kernel methods significantly enhance the performance of SVMs in handling nonlinear classification tasks; finding applications across machine learning, pattern recognition, and data mining [chandra2021survey].
Despite their advantages, kernel methods have limitations. They can be computationally expensive, particularly with large datasets, due to pairwise similarity calculations. Additionally, they risk overfitting when the dimensionality of the feature space increases. Selecting an appropriate kernel function requires substantial domain knowledge or extensive experimentation, and the resulting models are less interpretable compared to linear approaches. These challenges, along with sensitivity to hyperparameter tuning and scalability concerns, highlight the trade-offs when employing kernel methods in complex classification problems [hofmann2008kernel].
Therefore, a recent method, introduced by Dagher [dagher2008quadratic] and further refined by Luo et al. [luo2016soft], seeks to explicitly separate data using a quadratic classifier directly in the space of sample points rather than a transformed feature space:
$f_{W,b,c}(x)=\frac{1}{2}x^{T}Wx+b^{T}x+c=0.$
The soft margin version of this model imposes penalties for misclassification to accommodate potential noise and outliers is as follows:
$\min_{W,b,c,\xi}$ $\sum_{i=1}^{m}\|W{x_{i}}+b\|_{2}^{2}+\mu\sum_{i=1}^{m}\xi_{i}$ (QSVM)
s.t. $y_{i}(\frac{1}{2}x_{i}^{T}W{x_{i}}+b^{T}x_{i}+c)\geq 1-\xi_{i},\quad i=1,\dots,m,$
$W\in\mathbb{S}_{n},\,b\in\mathbb{R}^{n},\,c\in\mathbb{R},\,\xi\in\mathbb{R}^{m}_{+}.$
Notice that utilizing this Kernel-free quadratic surface idea for data separation exponentially increases the model’s degrees of freedom. Specifically, the parameter count in this model scales as $\mathcal{O}(n^{2})$, contrasting with traditional SVMs that possess parameters $\mathcal{O}(n)$. Although this complexity offers additional adaptability in training set segregation, it also increases the risk of overfitting, potentially compromising classification accuracy on testing datasets. This concern is particularly pronounced when dealing with linearly separable training data; in such cases, the expectation is for QSVMs to identify a suitable hyperplane for data division. However, QSVM does not guarantee this outcome. On the other hand, $\ell_{1}$ norm regularization techniques have demonstrated efficacy in reducing model complexity across various applications. In particular, [mousavi2022quadratic] advocates for the integration of $\ell_{1}$-regularization into QSVMs:
$\min_{W,b,c,\xi}$ $\sum_{i=1}^{m}\|W{x_{i}}+b\|_{2}^{2}+\lambda\sum_{1\leq i\leq j\leq n}|W_{ij}|+\mu\sum_{i=1}^{m}\xi_{i}$ (L1-QSVM)
s.t. $y_{i}(\frac{1}{2}x_{i}^{T}W{x_{i}}+b^{T}x_{i}+c)\geq 1-\xi_{i};\quad\forall i\in[m],$
$W\in\mathbb{S}_{n},\,b\in\mathbb{R}^{n},\,c\in\mathbb{R},\,\xi\in\mathbb{R}^{m}_{+},$
where $\mu>0$ is a positive penalty for incorporating noise and outliers. The authors demonstrate that the above model reduces to the standard SVM when $\lambda$ is sufficiently large. Other sparsity surrogates, such as the use of $\ell_{p}$ norms for $p\in(0,1)$, have also been explored in the literature on linear SVMs, feature selection, and K-support vector classification regression. However, the non-convexity of these models poses significant challenges for algorithm design [moosaei2024sparse, moosaei2022bounds, yao2017sparse, blanco2020lp, landeros2023algorithms, ma2019sparse].
On the other hand, it is known that $\ell_{0}$ directly enforces sparsity by minimizing the number of non-zero elements, unlike $\ell_{p}$ for $0<p\leq 1$, which only approximates sparsity and may introduce unnecessary regularization. For example, in compressive sensing, the $\ell_{0}$ norm achieves higher compression rates compared to its convex and nonconvex relaxations $\ell_{p}$ norms, making it a more effective tool for sparse representation. Therefore, we investigate $\ell_{0}$-norm based kernel-free quadratic SVMs and introduce a penalty decomposition algorithm for effectively tackling them in the next section.
Kernel-free quadratic surface models deviate from the conventional quadratic program formats in quadratic optimization. Hence, several notations and definitions must be introduced to rectify this and align them accordingly. We start with a square matrix $A=[a_{ij}]_{i=1,\dots,n;j=1,\dots,n}\in\mathbb{R}^{n\times n}$. Its vectorization is straightforwardly represented as $\mbox{vec}(A):=\left[a_{11},\dots,a_{n1},a_{12},\dots,a_{n2},\dots,a_{1n},\dots,a_{nn}\right]^{T}\in\mathbb{R}^{n^{2}}$. However, if $A$ is symmetric, $\mbox{vec}(A)$ carries redundant data, hence we opt for its half-vectorization:
$\mbox{hvec}(A):=\left[a_{11},\dots,a_{n1},a_{22},\dots,a_{2n},\dots,a_{nn}\right]^{T}\in\mathbb{R}^{\frac{n(n+1)}{2}}.$
For any $n\in\mathbb{N}$, there exists a unique elimination matrix $L_{n}\in\mathbb{R}^{\frac{n(n+1)}{2}\times n^{2}}$ such that $L_{n}\mbox{vec}(A)=\mbox{hvec}(A)$ for any $A\in S_{n}$, where $S_{n}$ denotes the set of symmetric matrices, as noted in [magnus1980elimination]. Additionally, this elimination matrix $L_{n}$ has a full row rank. Conversely, for any $n\in\mathbb{N}$, there exists a unique duplication matrix $D_{n}\in\mathbb{R}^{n^{2}\times\frac{n(n+1)}{2}}$ such that $D_{n}\mbox{hvec}(A)=\mbox{vec}(A)$ for any $A\in S_{n}$, and $L_{n}D_{n}=I_{\frac{n(n+1)}{2}}$.
Definition 2.1.
For $i\in[m],$ let
$s_{i}:=$ $\frac{1}{2}\mbox{hvec}(x_{i}x_{i}^{T}),\quad r_{i}:=[s_{i};x_{i}],\quad w:=\mbox{hvec}(W),\quad z:=[w;b],\quad V:=\begin{bmatrix}I_{\frac{n(n+1)}{2}}&0_{\frac{n(n+1)}{2}\times n}\end{bmatrix},$
$X_{i}:=$ $I_{n}\otimes x_{i}^{T},\quad M_{i}:=X_{i}D_{n},\quad H_{i}:=\begin{bmatrix}M_{i}&I_{n}\end{bmatrix},\quad G:=2\sum_{i=1}^{m}H_{i}^{T}H_{i},\quad X:=[x_{1},x_{2},\dots,x_{m}]^{T}.$
Thus, for fixed $i\in[m]$, we get the following equations:
$\left\{\begin{array}[]{ll}Wx_{i}=X_{i}\mbox{vec}(W)\,=X_{i}D_{n}\mbox{hvec}(W)\,=\,M_{i}\mbox{hvec}(W)\,=\,M_{i}w,&\\ Wx_{i}+b=\,M_{i}w+I_{n}b\,=\,H_{i}z,&\\ \frac{1}{2}x_{i}^{T}Wx_{i}+b^{T}x_{i}+c=z^{T}r_{i}+c,&\\ \sum_{i=1}^{m}\|Wx_{i}+b\|_{2}^{2}=\sum_{i=1}^{m}\left(H_{i}z\right)^{T}\left(H_{i}z\right)\,=\,z^{T}(\sum_{i=1}^{m}H_{i}^{T}H_{i})z=\frac{1}{2}z^{T}Gz.\end{array}\right.$
3 Sparse $\ell_{0}$-norm based Kernel-free Quadratic Surface SVMs and An effective Algorithm
It is empirically observed that real-world datasets are often governed by only a few key parameters, reinforcing the need for sparse $\ell_{0}$-norm based models. Such models are essential for ensuring better interpretability and practical applicability by identifying and leveraging these critical parameters. However, while $\ell_{p}$-norms ($p\in(0,1]$) are often used as surrogates for $\ell_{0}$, they approximate sparsity rather than enforcing it directly, which can result in solutions with near-zero values instead of exact zeros. Additionally, the sparsity level achieved with $\ell_{p}$-norms depends on parameter tuning and cannot be explicitly controlled, making them less effective in capturing true sparsity patterns. These limitations, along with their computational challenges, further highlight the importance of employing $\ell_{0}$-norm based approaches for tasks requiring precise sparsity.
In this paper, to mitigate the overfitting caused by the over-parameterization of the symmetric matrix $W$ and the vector $b$ in the quadratic classifier $f$, we impose a cardinality constraint on these parameters. This constraint effectively reduces the model’s complexity and enhances its generalization ability. Specifically, we study two sparse $\ell_{0}$-norm based, kernel-free quadratic surface SVM models. The first model employs the hinge loss function and is formulated as follows:
$\min_{W,b,c,\xi}$ $\sum_{i=1}^{m}\|W{x_{i}}+b\|_{2}^{2}+\mu\sum_{i=1}^{m}\xi_{i}$ (L0-QSVM)
s.t. $y_{i}(\frac{1}{2}x_{i}^{T}W{x_{i}}+b^{T}x_{i}+c)\geq 1-\xi_{i};\quad\forall i\in[m],$
$\|[\mbox{hvec}(W);b]\|_{0}\leq k,$
$W\in\mathbb{S}_{n},\,b\in\mathbb{R}^{n},\,c\in\mathbb{R},\,\xi\in\mathbb{R}^{m}_{+}.$
Using the quadratic loss function, we derive the least-squares version of the model as follows:
$\min_{W,b,c,\xi}$ $\sum_{i=1}^{m}\|Wx_{i}+b\|_{2}^{2}+\mu\sum_{i=1}^{m}\xi^{2}_{i}$ (LS-L0-QSVM)
s.t. $y_{i}(\frac{1}{2}x_{i}^{T}W{x_{i}}+b^{T}x_{i}+c)=1-\xi_{i};\quad\forall i\in[m],$
$\|[\mbox{hvec}(W);b]\|_{0}\leq r,$
$W\in\mathbb{S}_{n},\,b\in\mathbb{R}^{n},\,c\in\mathbb{R},\,\xi\in\mathbb{R}^{m},$
To present our penalty decomposition algorithm, we introduced the following unified model:
$\min_{W\in\mathbb{S}_{n},\,b\in\mathbb{R}^{n},\,c\in\mathbb{R}}$ $\sum_{i=1}^{m}\|Wx_{i}+b\|_{2}^{2}+\mu\sum_{i=1}^{m}\mathrm{H}(1-y_{i}f_{W,b,c}(x_{i}))\qquad\mbox{s.t.}\qquad\|[\mbox{hvec}(W);b]\|_{0}\leq k,$
which reduces to L0-QSVM for the well-received hinge loss $\mathrm{H}(t)=\max(t,0)$ and reduces to LS-L0-QSVM’ for the quadratic loss $\mathrm{H}(t)=t^{2}$. Based on the notation introduced in the previous section, the above model is equivalent to:
$\min_{z\in\mathbb{R}^{\frac{n(n+1)}{2}+n},c\in\mathbb{R}}$ $\frac{1}{2}z^{T}Gz+\mu\sum_{i=1}^{m}\mathrm{H}(1-y_{i}(z^{T}r_{i}+c))\qquad\mbox{s.t.}\qquad\|z\|_{0}\leq k.$ (1)
We now apply an effective penalty decomposition method for solving 1 as follows. After introducing a new variable $u$, this problem becomes:
$\min_{z,u\in\mathbb{R}^{\frac{n(n+1)}{2}+n},c\in\mathbb{R}}$ $\frac{1}{2}z^{T}Gz+\mu\sum_{i=1}^{m}\mathrm{H}(1-y_{i}(z^{T}r_{i}+c))\qquad\mbox{s.t.}\qquad\|z\|_{0}\leq k\qquad\mbox{and}\qquad z-u=0.$
Thus, by penalizing the last constraint above, the problem 1 can be addressed by solving a sequence of penalty subproblems as follows:
$\min_{z,u\in\mathbb{R}^{\frac{n(n+1)}{2}+n},c\in\mathbb{R}}$ $q_{\rho}(z,c,u):=\frac{1}{2}z^{T}Gz+\mu\sum_{i=1}^{m}\mathrm{H}(1-y_{i}(z^{T}r_{i}+c))+\frac{1}{2}\rho\|z-u\|_{2}^{2}\qquad\mbox{s.t.}\qquad\|u\|_{0}\leq k.$ ($P_{\rho}$)
Suppose we have a feasible point denoted as $(z,c)^{\text{feas}}$ of 1 in hand. Then, let:
$\Upsilon\geq\max\{f((z,c)^{\text{feas}}),\min_{z,c}q_{\rho^{(0)}}(z,c,u^{(0)}_{0})\}>0.$
We solve the above penalty subproblem for a fixed positive $\rho$ via a block coordinate descent method. Therefore, we have the following sparse $\ell_{0}$-norm based penalty decomposition for solving (1).
Algorithm 1 Sparse Quadratic Surface Penalty Decomposition (SQS-PD)
1:  Inputs: $\rho^{(0)}>0,\beta>1,k$, and $u^{(0)}_{0}$ such that $\|u^{(0)}_{0}\|_{0}\leq k$.
2:  $j\leftarrow 0$.
3:  repeat
4:     $l\leftarrow 0$.
5:     repeat
6:         $(z^{(j)}_{l+1},c^{(j)}_{l+1})\leftarrow\mbox{Argmin}_{(z,c)\in\mathbb{R}^{\frac{n(n+1)}{2}+n}\times\mathbb{R}}\ q_{\rho^{(j)}}(z,c,u^{(j)}_{l}).$
7:         $u^{(j)}_{l+1}\leftarrow\mbox{Argmin}_{\{u|\|u\|_{0}\leq k\}}\ q_{\rho^{(j)}}(z^{(j)}_{l+1},c^{(j)}_{l+1},u).$
8:        $l\leftarrow l+1$.
9:     until stopping criterion (2) is met.
10:      $(z^{(j)},c^{(j)},u^{(j)})\leftarrow(z^{(j)}_{l},c^{(j)}_{l},u^{(j)}_{l})$.
11:     $\rho^{(j+1)}\leftarrow\beta\cdot\rho^{(j)}$.
12:     If $\min_{(z,c)\in\mathbb{R}^{\frac{n(n+1)}{2}+n}\times\mathbb{R}}\ q_{\rho^{(j+1)}}(z,c,u^{(j)})>\Upsilon$, then $u_{0}^{(j+1)}\leftarrow z^{\text{feas}}$. Otherwise, $u^{(j+1)}_{0}\leftarrow u^{(j)}_{l}$.
13:     $j\leftarrow j+1$.
14:  until stopping criterion (3) is met.
The stopping criterion for the inner loop is the following:
$\max\left\{\frac{\|z_{l}-z_{l-1}\|_{\infty}}{\max\left(\|z_{l}\|_{\infty},1\right)},\frac{\|c_{l}-c_{l-1}\|_{\infty}}{\max\left(\|c_{l}\|_{\infty},1\right)},\frac{\|u_{l}-u_{l-1}\|_{\infty}}{\max\left(\|u_{l}\|_{\infty},1\right)},\right\}\leq\epsilon_{I}.$ (2)
and for the outer loop is:
$\|z^{(j)}-u^{(j)}\|_{\infty}\leq\epsilon_{O}.$ (3)
According to Theorem 4.3 in [lu2013sparse] and Theorem 3.34 in [ruszczynski2011nonlinear] Algorithm 1 obtains a first-order optimality point of 1. Therefore, our focus will be next on effectively solving the restricted subproblems within this algorithm.
Observe that the associated subproblem for $u$ is:
$\min_{u\in\mathbb{R}^{\frac{n(n+1)}{2}+n}}\ \|z-u\|_{2}^{2}\qquad\mbox{s.t.}\qquad\|u\|_{0}\leq k;$ (4)
which obtains a solution
$u=[z_{\mathcal{J}};0],$ (5)
where $\mathcal{J}$ contains indices of $k$ largest components of $z$ in absolute value, that is, Step 7 of Algorithm 1 has a closed-form solution. Next, we start discussing how to tackle the associated subproblem to $z$ for each of the loss functions.
3.1 Sparse $\ell_{0}$-norm based Kernel-free Quadratic Surface SVM Model
For the hinge loss function $\mathrm{H}(t)=\max(0,t)$, the sparse quadratic model 1 reduces to L0-QSVM, which after using the notation in Section 2 is equivalent to:
$\min_{z,c,\xi}$ $\frac{1}{2}{z}^{T}Gz+\mu\sum_{i=1}^{m}\xi_{i}$ (L0-QSVM’)
s.t. $y_{i}(z^{T}r_{i}+c)\geq 1-\xi_{i};\quad\forall i\in[m],$
$\|z\|_{0}\leq r.$
$z\in\mathbb{R}^{\frac{n(n+1)}{2}+n},c\in\mathbb{R},\xi\in\mathbb{R}^{m}_{+}.$
Recalling $P_{\rho}$, the subproblem associated to $z$ for the hinge loss function becomes:
$\min_{z\in\mathbb{R}^{\frac{n(n+1)}{2}+n},c\in\mathbb{R}}$ $\frac{1}{2}z^{T}Gz+\mu\sum_{i=1}^{m}\max(1-y_{i}(z^{T}r_{i}+c),0)+\frac{1}{2}\rho\|z-u\|_{2}^{2},$
which is equivalent to the following feasible and bounded below the convex quadratic optimization problem:
$\min_{z,c,\xi}$ $\frac{1}{2}{z}^{T}(G+\rho I)z-\rho u^{T}z+\mu\sum_{i=1}^{m}\xi_{i}$ ($P^{h}_{\rho,(z,c,\xi)}$)
s.t. $y_{i}(z^{T}r_{i}+c)\geq 1-\xi_{i};\quad\forall i\in[m]$
$z\in\mathbb{R}^{\frac{n(n+1)}{2}+n},c\in\mathbb{R},\xi\in\mathbb{R}^{m}_{+},.$
Even though a standard solver can be used for this problem, standard practice in the SVM literature encourages applying the strong duality theorem in convex programming. Thus, let us bring the Lagrangian function of this model:
$L(z,c,\xi,\alpha)$ $=$ $\frac{1}{2}{z}^{T}(G+\rho I)z-\rho u^{T}z+\mu\sum_{i=1}^{m}\xi_{i}-\sum_{i=1}^{m}\alpha_{i}(y_{i}(z^{T}r_{i}+c)-1+\xi_{i})$
$=$ $\frac{1}{2}{z}^{T}(G+\rho I)z-z^{T}(\sum_{i=1}^{m}\alpha_{i}y_{i}r_{i}+\rho u)+\sum_{i=1}^{m}(\mu-\alpha_{i})\xi_{i}-c\sum_{i=1}^{m}\alpha_{i}y_{i}.$
Therefore, the dual problem of $P^{h}_{\rho,(z,c,\xi)}$ becomes:
$\min_{\alpha\in\mathbb{R}^{m}}$ $-\sum_{i=1}^{m}\alpha_{i}+\frac{1}{2}\left(\sum_{i=1}^{m}y_{i}\alpha_{i}r_{i}+\rho u\right)^{T}(G+\rho I)^{-1}\left(\sum_{j=1}^{m}y_{j}\alpha_{j}r_{j}+\rho u\right)$ ($D_{\alpha}$)
s.t. $\sum_{i=1}^{m}y_{i}\alpha_{i}=0\qquad\mbox{and}\qquad 0\leq\alpha\leq\textbf{1}\mu.$
Therefore, the KKT conditions are:
$\left\{\begin{array}[]{ll}z=(G+\rho I)^{-1}(\sum_{i=1}^{m}\alpha_{i}y_{i}r_{i}+\rho u)&\\ \alpha_{i}(1-\xi_{i}-y_{i}(z^{T}r_{i}+c))=0;\quad\forall i\in[m]&\\ (\mu-\alpha_{i})\xi_{i}=0;\quad\forall i\in[m]&\\ y_{i}(z^{T}r_{i}+c)\geq 1-\xi_{i};\quad\forall i\in[m]&\\ \sum_{i=1}^{m}y_{i}\alpha_{i}=0;\quad\forall i\in[m]&\\ \xi\in\mathbb{R}^{m}_{+}\qquad\mbox{and}\qquad 0\leq\alpha\leq\textbf{1}\mu,&\\ \end{array}\right.$
which implies that, once a solution to the dual $D_{\alpha}$ is provided, a solution to $P^{h}_{\rho,(z,c,\xi)}$ can be obtained from:
$z=(G+\rho I)^{-1}(\sum_{\alpha_{i}\neq 0}\alpha_{i}y_{i}r_{i}+\rho u),\qquad c=\max_{y_{i}=1,\alpha_{i}>0}-z^{T}r_{i},\qquad\xi_{i}=\left\{\begin{array}[]{ll}1-y_{i}(z^{T}r_{i}+c)\quad\mbox{if}\quad\alpha_{i}>0,&\\ 0\quad\mbox{if}\quad\alpha_{i}=0.&\\ \end{array}\right.$ (6)
Corollary 3.1.
For the hinge loss function, step 6 of Algorithm 1 can be obtained by solving $D_{\alpha}$ and then applying 6.
3.2 Least-squares Sparse $\ell_{0}$-norm based Kernel-free Quadratic Surface SVM Model
For the quadratic loss function $\mathrm{H}(t)=t^{2}$, the sparse quadratic model 1 reduces to the following least-squares sparse $\ell_{0}$-norm based kernel-free quadratic surface SVM model LS-L0-QSVM, which through the notations presented in Section 2 becomes:
$\min_{z,c,\xi}$ $\frac{1}{2}{z}^{T}Gz+\mu\sum_{i=1}^{m}\xi^{2}_{i}$ (LS-L0-QSVM’)
s.t. $y_{i}(z^{T}r_{i}+c)=1-\xi_{i};\quad\forall i\in[m],$
$\|z\|_{0}\leq r,$
$z\in\mathbb{R}^{\frac{n(n+1)}{2}+n},c\in\mathbb{R},\xi\in\mathbb{R}^{m},.$
The associated subproblem to $z$ in this case is equivalent to:
$\min_{z\in\mathbb{R}^{\frac{n(n+1)}{2}+n},c\in\mathbb{R}}$ $\frac{1}{2}{z}^{T}Gz+\mu\sum_{i=1}^{m}(1-y_{i}(z^{T}r_{i}+c))^{2}+\frac{1}{2}\rho\|z-u\|^{2}_{2}.$ ($P_{(z,c)}^{q}$)
By letting $A=[r_{1},r_{2},\dots,r_{m}]^{T}$ and $D={\rm{\mbox{\bf{diag}}}}(y_{1},y_{2},\dots,y_{m})$, the objective function of the latter is equal to $T=\frac{1}{2}z^{T}(G+\rho I)z+\mu\|\textbf{1}-D(Az+c\textbf{1})\|_{2}^{2}-\rho u^{T}z+\rho\|u\|_{2}^{2}$. Hence,
$\frac{\partial T}{\partial z}$ $=$ $(G+\rho I)z-2\mu A^{T}D(\textbf{1}-D(Az+c\textbf{1}))-\rho u$
$\frac{\partial T}{\partial c}$ $=$ $-2\mu\textbf{1}^{T}D(\textbf{1}-D(Az+c\textbf{1}))$
which can be integrated into the following:
$\begin{bmatrix}G+\rho I+2\mu A^{T}D^{2}A&2\mu A^{T}D\textbf{1}\\ 2\mu\textbf{1}^{T}D^{2}A&2\mu\textbf{1}^{T}D^{2}\textbf{1}\end{bmatrix}\begin{bmatrix}z\\ c\end{bmatrix}=\begin{bmatrix}2\mu A^{T}D\textbf{1}+\rho u\\ 2\mu\textbf{1}^{T}D\textbf{1},\end{bmatrix}$
which has a positive definite coefficient matrix such that:
$\begin{bmatrix}z\\ c\end{bmatrix}=\begin{bmatrix}G+\rho I+2\mu A^{T}D^{2}A&2\mu A^{T}D\textbf{1}\\ 2\mu\textbf{1}^{T}D^{2}A&2\mu\textbf{1}^{T}D^{2}\textbf{1}\end{bmatrix}^{-1}\begin{bmatrix}2\mu A^{T}D\textbf{1}+\rho u\\ 2\mu\textbf{1}^{T}D\textbf{1},\end{bmatrix}$ (7)
Corollary 3.2.
For the quadratic loss function, step 6 of Algorithm 1 is obtained from (7).
To conclude this section, we introduced two sparse $\ell_{0}$-norm based kernel-free quadratic surface SVM models, utilizing hinge and quadratic loss functions, respectively. Due to the cardinality constraint in these models, they are NP-hard, necessitating an efficient algorithm for practical implementation. To address this, we unified the models under a general framework 1 and proposed a penalty decomposition algorithm that introduces an auxiliary variable $u$. The subproblem (4) associated with $u$ admits a straightforward closed-form solution (5). Specifically, for the hinge-loss based model L0-QSVM, we demonstrated that the subproblem for $z$ is best addressed through convex duality. Solving the dual problem yields the primal solution via (6). On the other hand, for the quadratic-loss based model LS-L0-QSVM, we showed that the subproblem for $z$ has a closed-form solution directly provided in (7). Thus, our iterative algorithm is both computationally efficient and effective.
4 Numerical Experiments
In this section, we conduct numerical experiments on real datasets to evaluate the performance of our proposed models L0-QSVM and LS-L0-QSVM when solved by Algorithm 1. We compare its accuracy and ability to detect true underlying sparsity patterns against the benchmark method L1-QSVM. To compare the performance of our proposed models L0-QSVM and LS-L0-QSVM with the standard SVM, quadratic kernel SVM, QSVM, and L1-QSVM, we implemented Algorithm 1 in Python. For linear and quadratic kernel SVMs, we utilized the Scikit-learn library. We evaluated the accuracy of these models on several balanced datasets from the UCI repository, including Iris, Glass Identification, Raisin, Wine Quality, Hepatitis C Virus for Egyptian patients (HCV), Yeast, and Wholesale Customers. These datasets are widely used benchmarks for evaluating the performance of newly developed models in the literature. They encompass both binary and multi-class label variables. For multi-class datasets, we utilized the one-vs-rest strategy combined with majority voting to determine the predicted class.
We performed hyperparameter tuning using random search. For the standard SVM, we optimized the regularization parameter $C$ by searching over values on a logarithmic scale from $10^{-6}$ to $10^{3}$. For the quadratic kernel SVM, we additionally tuned the parameters $\gamma$ and coef0 over a logarithmic scale from $10^{-6}$ to $10$. For L1-QSVM and L0-QSVM and LS-L0-QSVM, which are kernel-free, we tuned $\mu$ on a logarithmic scale from $10^{-2}$ to $100$. For L0-QSVM and LS-L0-QSVM, we also tuned $k$, the number of selected features, over the range $1$ to $2^{\text{number of features}}$, and compared performance using both hinge loss and quadratic loss.
To evaluate model performance, we employed 5-fold cross-validation. The dataset was partitioned into five folds, with each fold serving as the test set in turn, while the remaining four folds were further split into three for training and one for validation. Testing was conducted using models tuned with the best hyperparameters selected from the validation set. For each fold, we sampled 100 different hyperparameter configurations for tuning. This process produced five accuracy values for each dataset, from which we calculated the mean and standard deviation, as reported in Table 1.
Table 1: Test Accuracy Comparisons
Dataset SVM Quad. Kernel QSVM L1-QSVM L0-QSVM LS-L0-QSVM
Iris 0.9533 ± 0.034 0.940 ± 0.057 0.953 ± 0.0618 0.953 ± 0.077 0.953 ± 0.034 0.960 ± 0.049
Glass 0.617 ± 0.049 0.692 ± 0.076 0.631 ± 0.041 0.640 ± 0.033 0.650 ± 0.051 0.656 ± 0.062
Raisin 0.870 ± 0.033 0.868 ± 0.029 0.862 ± 0.030 0.863 ± 0.029 0.859 ± 0.030 0.851 ± 0.031
Wine 0.961 ± 0.030 0.955 ± 0.038 0.955 ± 0.022 0.961 ± 0.029 0.977 ± 0.021 0.961 ± 0.034
HCV 0.951 ± 0.011 0.951 ± 0.010 0.936 ± 0.013 0.939 ± 0.013 0.941 ± 0.005 0.924 ± 0.019
Yeast 0.586 ± 0.012 0.591 ± 0.016 0.610 ± 0.024 0.604 ± 0.030 0.590 ± 0.027 0.567 ± 0.028
Wholesale 0.907 ± 0.024 0.896 ± 0.043 0.902 ± 0.023 0.905 ± 0.027 0.900 ± 0.022 0.893 ± 0.016
Next, to illustrate the strength of the $\ell_{0}$ norm in identifying effective sparse patterns that perform well in practice, we visualize these patterns in both 2D and 3D. Figure 1 demonstrates that L0-QSVM retains high performance even in lower-dimensional spaces that are visualizable. The figure shows the decision boundaries of L0-QSVM for the Iris dataset with $k=2$. In Figure 2, the Wine Quality dataset, the decision boundary appears as a hyperplane separating the classes, whereas in the Raisin dataset, it is depicted as a curved 3D manifold enclosing one class and excluding another. These examples highlight the potential of sparse $\ell_{0}$ quadratic surface SVMs for data visualization, enabling projection into two- or three-dimensional feature spaces.
Figures 3, 4, and 5 illustrate a comparison of the weight matrices $W$ and vectors $b$ for L1-QSVM and L0-QSVM when applied to the Iris, Glass, and Heart datasets. The weight matrices were obtained using the optimal hyperparameters for each model, determined through validation on a randomly selected 20% subset of the Iris, Glass, and Heart datasets. As evident from the figures, the $W$ matrices and $b$ vectors for L0-QSVM exhibit significantly greater sparsity compared to those of L1-QSVM. This sparsity substantially enhances the interpretability of our classifiers, surpassing what is suggested by their $\ell_{1}$-norm counterparts. Nevertheless, despite this increased sparsity, Table 1 shows that the accuracy of L0-QSVM is either higher or comparable to that of QSVM.
Figure 1: L0-QSVM Classifier for $k=2$ on Iris dataset
Figure 2: Surfaces obtained by L0-QSVM with $k=3$ on Raisin and Wine datasets
Figure 3: Magnitudes of optimal parameters (weight matrix $W$ and vector $b$) of classifiers found by L1-QSVM (a) and L0-QSVM (b) for the Iris dataset.
Figure 4: Magnitudes of optimal parameters (weight matrix $W$ and vector $b$) of classifiers found by L1-QSVM (a) and L0-QSVM (b) for the Glass dataset.
Figure 5: Magnitudes of optimal parameters (weight matrix $W$ and vector $b$) of classifiers found by L1-QSVM (a) and L0-QSVM (b) for the Heart dataset.
Figure 6 illustrates the accuracy of L0-QSVM across different values of $k$. As observed, these models often achieve their full predictive power with relatively small values of $k$, while maintaining the flexibility of models incorporating second-order features. Unlike traditional kernel methods that transform the data into a higher-dimensional feature space, L0-QSVM operates directly in the original feature space, enhancing interpretability. This quality is particularly valuable in domains like healthcare, where model interpretability is a critical requirement.
Figure 6: Accuracy vs sparsity level $k$
5 Conclusion
This paper introduces sparse $\ell_{0}$-norm based kernel-free quadratic surface SVM models to address challenges such as overfitting, interpretability, and the complexity of kernel-free quadratic surface machines in binary classification tasks. By incorporating a cardinality constraint on model parameters, these methods effectively reduce over-parameterization while enhancing generalization and sparsity. However, these models are inherently intractable and NP-hard, underscoring the need for an effective algorithm to solve them. To this end, we developed a penalty decomposition algorithm that leverages duality theory and provides closed-form solutions for key subproblems. Numerical experiments conducted on real-world datasets demonstrate that the proposed models outperform standard linear and quadratic SVM models and $\ell_{1}$-regularized counterparts in detecting true underlying sparse patterns while maintaining high classification accuracy. These results underscore the practical advantages of the $\ell_{0}$-norm based approach in applications requiring interpretable and sparse decision boundaries.
The proposed framework holds significant promise for fields where transparency and interpretability are critical, such as healthcare and finance. In addition to its standard formulation, the twin version of the proposed models enhances efficiency by solving smaller, decoupled subproblems, making it suitable for high-dimensional data. Moreover, the framework’s ability to handle imbalance scenarios is a notable strength, as it effectively incorporates sparsity constraints to mitigate the dominance of majority classes, ensuring better representation of minority classes. Future research could explore scalability to larger datasets, adaptive methods for hyperparameter tuning, extensions to multi-class classification problems, strategies to improve twin formulations, and techniques to address imbalance in more complex real-world datasets, thereby further enhancing the applicability and robustness of the models across diverse settings.
[…]