### v3

Abstract
Kernel-free quadratic surface support vector machines have recently gained traction due to their flexibility in modeling nonlinear decision boundaries without relying on kernel functions. However, the introduction of a full quadratic classifier significantly increases the number of model parameters, scaling quadratically with data dimensionality, which often leads to overfitting and makes interpretation difficult. To address these challenges, we propose a sparse variant of the QSVM by enforcing a cardinality constraint on the model parameters. While enhancing generalization and promoting sparsity, leveraging the $\ell_{0}$-norm inevitably incurs additional computational complexity. To tackle this, we develop a penalty decomposition algorithm capable of producing solutions that provably satisfy the first-order Lu-Zhang optimality conditions. Our approach accommodates both hinge and quadratic loss functions. In both cases, we demonstrate that the subproblems arising within the algorithm either admit closed-form solutions or can be solved efficiently through dual formulations, which contributes to the method’s overall effectiveness. We also analyze the convergence behavior of the algorithm under both loss settings. Finally, we validate our approach on several real-world datasets, demonstrating its ability to reduce overfitting while maintaining strong classification performance. The complete implementation and experimental code are publicly available at: https://github.com/raminzandvakili/L0-QSVM.
keywords: binary classification, quadratic support vector machines, $\ell_{0}$ norm, penalty decomposition method
1 Introduction
Soft-margin support vector machines (SVMs) and their extensions remain central to modern binary classification, as they aim to find decision boundaries that maximize margin while minimizing misclassification [1, 2, 3]. In parallel, sparse linear models have become increasingly important due to their interpretability and reduced risk of overfitting, achieved by explicitly limiting model complexity [4, 5, 6, 7, 8]. However, linear sparsity alone cannot capture nonlinear structures in complex data. Kernel-based SVMs address this by mapping inputs to high-dimensional feature spaces via kernel functions. While effective, this comes at a cost: the resulting models often lack transparency and depend heavily on kernel and hyperparameter choices, making interpretation difficult and computational costs high.
To maintain interpretability without resorting to kernelization, kernel-free quadratic surface support vector machines (QSVMs) have been proposed. These models directly learn quadratic separating surfaces and optimize approximate geometric margins while controlling for noise and outliers [9]. Additionally, their explicit functional form allows the inclusion of prior information, such as Universum data, and variants like twin QSVMs improve scalability by decomposing the original task into smaller subproblems [10]. Yet, a key drawback of these models is that their number of parameters scales quadratically with the input dimension, i.e., $\mathcal{O}(n^{2})$, making them prone to overfitting and limiting their utility on even moderately sized datasets.
A recent model that constrains the quadratic weight matrix to be diagonal [11] reduces the parameter space to $\mathcal{O}(n)$ and helps mitigate overfitting, but it implicitly assumes negligible feature interactions—an assumption rarely valid in practice—and risks discarding informative pairwise dependencies, leading to underfitting. More flexible approaches employ regularization strategies with different trade-offs: $\ell_{1}$-norm methods [12, 10, 13] promote sparsity but yield nonunique solutions; $\ell_{2}$-norm regularization [14] controls overfitting without inducing sparsity; and $\ell_{p}$-norm formulations [15] offer tunable sparsity but require more careful parameter selection.
Sparse recovery theory establishes that $p>1$ typically yields dense solutions [16], $p=1$ offers convexity but weaker sparsity than $0<p<1$, and within the latter range, $p\in[1/2,1)$ often delivers the sparsest solutions, with smaller $p$ providing marginal additional benefit [17]. In contrast, $\ell_{0}$-norm regularization uniquely enforces exact control over the number of nonzero parameters—a capability unattainable with any $\ell_{p}$ surrogate for p>0p>0—thereby achieving exact sparsity and offering stronger theoretical recovery guarantees. With recent advances in optimization, including penalty decomposition [18], greedy algorithms [19], and branch-and-bound methods [20], solving $\ell_{0}$-regularized models directly has become practical, making reliance on surrogate formulations increasingly unnecessary and, in many settings, suboptimal.
Building on these insights, we develop a novel class of kernel-free QSVMs that employ $\ell_{0}$ regularization to achieve exact sparsity in both the weight matrix and linear term. Our method combines the representational capacity of quadratic decision boundaries with the simplicity of sparse models, offering both powerful nonlinear classification and clear interpretability. The $\ell_{0}$ constraint provides solutions with precisely $k$ non-zero elements, giving direct control over model complexity while automatically performing feature selection. Unlike alternative approaches such as the two-level $\ell_{1}$-norm [21] or sparse group $k$-max regularization [22] that only approximate sparsity through soft thresholding, our method enforces exact sparsity by design. These alternatives require careful tuning of additional parameters (like the top-$k$ support size or shrinkage weights $\rho$) and can leave residual small coefficients. Our $\ell_{0}$-based formulation eliminates these heuristic choices while guaranteeing truly sparse solutions.
To address the computational challenges posed by the inclusion of an $\ell_{0}$ regularization term, we develop a penalty decomposition algorithm that recasts the original problem into a sequence of more tractable subproblems, drawing inspiration from the general framework in [23, 18]. By introducing an auxiliary variable that decouples the $\ell_{0}$-norm constraint, our method alternates between updates that are either analytically tractable or efficiently solvable using duality. In particular, for the hinge loss setting, each iteration reduces to solving a convex QSVM using dual optimization, while in the quadratic loss case, the subproblems admit closed-form updates via linear algebra. These properties make the method both scalable and precise in controlling sparsity.
The remainder of the paper is structured as follows. Section 2 reviews relevant literature. Section 3 presents our $\ell_{0}$-regularized QSVM models and describes the decomposition algorithm. Section 4 provides empirical results, and Section 5 concludes with a discussion of future directions.
2 Preliminaries and Related Work
This section briefly reviews several quadratic surface support-vector machine (QSVM) models. Consider a binary classification task with a training set
$\mathcal{D}=\{\left({x_{i}},y_{i}\right)_{i=1,\dots,m}\ \left|\ {x_{i}}\,\in\,\mathbb{R}^{n},\ y_{i}\in\{-1,1\}\right$.\},
where $m$ is the sample size, $n$ is the number of features, and $y_{i}$ is the label of $x_{i}$. $\mathcal{D}$ is called quadratically separable [9] if there exist $W\in\mathbb{R}^{n\times n},b\in\mathbb{R}^{n},\,\text{and}\,c\in\mathbb{R}$ if $y_{i}(x_{i}^{\top}W{x_{i}}+b^{\top}{x_{i}}+c)>0$ for every $i\in[m]$. A linearly separable dataset is a special case of quadratic separability obtained by setting $W=0$.
Standard SVMs seek a hyperplane that approximately separates a training set. In many real-world binary classification problems, however, data exhibit non-linear patterns, making linear separation difficult. Kernel methods overcome this limitation by mapping the data into a higher-dimensional feature space in which non-linear relationships appear linear. In that space, SVMs can identify optimal hyperplanes that more effectively capture complex structures. Consequently, kernel techniques markedly improve SVM performance on non-linear classification tasks and are widely applied in machine learning, pattern recognition, and data mining [24].
Despite their advantages, kernel methods also have notable limitations. They are computationally expensive because pair-wise similarity calculations scale poorly with large data sets. Moreover, as the dimensionality of the feature space grows, kernel models become prone to overfitting. Choosing a kernel function demands domain expertise or extensive experimentation, and the resulting classifiers are less interpretable than their linear counterparts. Sensitivity to hyper-parameter tuning and limited scalability further underscore the trade-offs associated with kernel methods in complex classification tasks [25].
Therefore, a recent method, introduced by Dagher [26] and further refined by Luo et al. [9], seeks to explicitly separate data using a quadratic classifier directly in the original input space, avoiding any kernel-induced feature mapping:
$f_{W,b,c}(x)=\frac{1}{2}x^{\top}Wx+b^{\top}x+c=0.$
The soft-margin variant of this model, which penalizes misclassifications to accommodate noise and outliers, is formulated as follows:
$\min_{W,b,c,\xi}$ $\sum_{i=1}^{m}\|W{x_{i}}+b\|_{2}^{2}+{\mathcal{C}}\sum_{i=1}^{m}\xi_{i}$ (QSVM)
s.t. $y_{i}(\frac{1}{2}x_{i}^{\top}W{x_{i}}+b^{\top}x_{i}+c)\geq 1-\xi_{i},\quad i=1,\dots,m,$
$W\in\mathbb{S}_{n},\,b\in\mathbb{R}^{n},\,c\in\mathbb{R},\,\xi\in\mathbb{R}^{m}_{+}.$
Notice that adopting a kernel-free quadratic surface framework dramatically enlarges the model’s degrees of freedom: the number of parameters grows on the order of $\mathcal{O}(n^{2})$, compared with the $\mathcal{O}(n)$ parameters in a standard linear SVM. While this added flexibility can improve the fit to the training data, it also raises the risk of overfitting and may diminish generalization to unseen samples. The danger is especially acute for linearly separable datasets, where one would prefer the QSVM to revert to a linear separator—yet QSVM offers no such guarantee. A widely used remedy is sparsity-promoting $\ell_{1}$-norm regularization; for example, Mousavi et al. [12] embed an $\ell_{1}$ penalty directly into the QSVM objective:
$\min_{W,b,c,\xi}$ $\sum_{i=1}^{m}\|W{x_{i}}+b\|_{2}^{2}+{\mathcal{C}}_{1}\sum_{1\leq i\leq j\leq n}|W_{ij}|+{\mathcal{C}}_{2}\sum_{i=1}^{m}\xi_{i}$ ($\ell_{1}$-QSVM)
s.t. $y_{i}(\frac{1}{2}x_{i}^{\top}W{x_{i}}+b^{\top}x_{i}+c)\geq 1-\xi_{i};\quad\forall i\in[m],$
$W\in\mathbb{S}_{n},\,b\in\mathbb{R}^{n},\,c\in\mathbb{R},\,\xi\in\mathbb{R}^{m}_{+},$
where ${\mathcal{C}}>0$ is a positive penalty for incorporating noise and outliers. The authors demonstrate that the above model reduces to the standard SVM when $\lambda$ is sufficiently large. Other $\ell_{0}$-norm surrogates, such as $\ell_{p}$ norms, for $p\in(0,1)$, have also been explored in the literature on linear SVMs, feature selection, and $K$-support vector classification regression. However, the non-convexity of these models poses significant challenges for algorithm design [27, 28, 29].
Although $\ell_{p}$ norms with $0<p\leq 1$ promote sparsity, only the $\ell_{0}$ norm enforces it directly by minimizing the number of non-zero parameters. In compressive sensing, for example, $\ell_{0}$ achieves higher compression rates than either the convex ($p=1$) or non-convex ($0<p<1$) $\ell_{p}$-norm relaxations, making it a superior tool for sparse representation. Motivated by this property, we next study sparse $\ell_{0}$-norm, kernel-free quadratic surface SVMs and present a penalty decomposition algorithm that solves these models efficiently.
Kernel-free quadratic surface models deviate from the conventional quadratic program formats in quadratic optimization. Hence, several notations and definitions must be introduced to rectify this and align them accordingly. We start with a square matrix $A=[a_{ij}]_{i=1,\dots,n;j=1,\dots,n}\in\mathbb{R}^{n\times n}$. Its vectorization is straightforwardly represented as $\mbox{vec}(A):=\left[a_{11},\dots,a_{n1},a_{12},\dots,a_{n2},\dots,a_{1n},\dots,a_{nn}\right]^{\top}\in\mathbb{R}^{n^{2}}$. However, if $A$ is symmetric, $\mbox{vec}(A)$ carries redundant data, hence we opt for its half-vectorization:
$\mbox{hvec}(A):=\left[a_{11},\dots,a_{n1},a_{22},\dots,a_{2n},\dots,a_{nn}\right]^{\top}\in\mathbb{R}^{\frac{n(n+1)}{2}}.$
For any $n\in\mathbb{N}$, there exists a unique elimination matrix $L_{n}\in\mathbb{R}^{\frac{n(n+1)}{2}\times n^{2}}$ such that $L_{n}\mbox{vec}(A)=\mbox{hvec}(A)$ for any $A\in S_{n}$, where $S_{n}$ denotes the set of symmetric matrices. Additionally, this elimination matrix $L_{n}$ has a full row rank. Conversely, for any $n\in\mathbb{N}$, there exists a unique duplication matrix $D_{n}\in\mathbb{R}^{n^{2}\times\frac{n(n+1)}{2}}$ such that $D_{n}\mbox{hvec}(A)=\mbox{vec}(A)$ for any $A\in S_{n}$, and $L_{n}D_{n}=I_{\frac{n(n+1)}{2}}$.
Definition 2.1.
For $i\in[m],$ let
$s_{i}:=$ $\frac{1}{2}\mbox{hvec}(x_{i}x_{i}^{\top}),\quad r_{i}:=[s_{i};x_{i}],\quad w:=\mbox{hvec}(W),\quad z:=[w;b],\quad V:=\begin{bmatrix}I_{\frac{n(n+1)}{2}}&0_{\frac{n(n+1)}{2}\times n}\end{bmatrix}$,
$X_{i}:=$ $I_{n}\otimes x_{i}^{\top},\quad M_{i}:=X_{i}D_{n},\quad H_{i}:=\begin{bmatrix}M_{i}&I_{n}\end{bmatrix},\quad G:=2\sum_{i=1}^{m}H_{i}^{\top}H_{i},\quad X:=[x_{1},x_{2},\dots,x_{m}]^{\top}.$
Thus, simply $G\succeq 0$. And, for fixed $i\in[m]$, we have the following equations:
$\left\{\begin{array}[]{ll}Wx_{i}=X_{i}\mbox{vec}(W)\,=X_{i}D_{n}\mbox{hvec}(W)\,=\,M_{i}\mbox{hvec}(W)\,=\,M_{i}w,&\\ Wx_{i}+b=\,M_{i}w+I_{n}b\,=\,H_{i}z,&\\ \frac{1}{2}x_{i}^{\top}Wx_{i}+b^{\top}x_{i}+c=z^{\top}r_{i}+c.\end{array}\right$.
Thus,
$\sum_{i=1}^{m}\|Wx_{i}+b\|_{2}^{2}=\sum_{i=1}^{m}\left(H_{i}z\right)^{\top}\left(H_{i}z\right)\,=\,z^{\top}\big{(}\sum_{i=1}^{m}H_{i}^{\top}H_{i}\big{)}z=\frac{1}{2}z^{\top}Gz.$
3 $\ell_{0}$-Regularized Quadratic Surface SVMs and A Penalty Decompotion Algorithm
Empirical evidence suggests that many real-world datasets, being governed by underlying physical laws, are effectively controlled by only a few dominant factors. This makes $\ell_{0}$-regularized models particularly appealing for capturing their essential structure. By isolating these critical components, $\ell_{0}$ regularization delivers greater interpretability and practical utility. Conversely, $\ell_{p}$ surrogates with $0<p\leq 1$ merely approximate sparsity: they shrink coefficients toward—but rarely all the way to—zero, and the degree of sparsity is controlled only indirectly through a tuning parameter. Because this parameter cannot set the exact number of non-zero entries, $\ell_{p}$ regularization often fails to capture the true zero pattern. These shortcomings highlight the benefits of employing $\ell_{0}$ regularization when precise sparsity is crucial. However, the inclusion of a sparsity constraint renders the model intractable, necessitating the development of an efficient solution method. To address this, we propose a penalty decomposition algorithm and demonstrate how its subproblems can be solved effectively. We will also study the convergence of this algorithm.
To mitigate overfitting arising from the over-parameterization of the symmetric matrix $W$ and vector $b$ in the quadratic classifier $f$, we impose sparsity constraints on these parameters. This restriction lowers model complexity and improves generalization. Accordingly, we examine two sparse $\ell_{0}$-norm kernel-free quadratic surface SVM formulations. The first employs the hinge loss and is expressed as follows:
$\min_{W,b,c,\xi}$ $\sum_{i=1}^{m}\|W{x_{i}}+b\|_{2}^{2}+{\mathcal{C}}\sum_{i=1}^{m}\xi_{i}$ ($\ell_{0}$-QSVM)
s.t. $y_{i}(\frac{1}{2}x_{i}^{\top}W{x_{i}}+b^{\top}x_{i}+c)\geq 1-\xi_{i};\quad\forall i\in[m],$
$\|[\mbox{hvec}(W);b]\|_{0}\leq k,$
$W\in\mathbb{S}_{n},\,b\in\mathbb{R}^{n},\,c\in\mathbb{R},\,\xi\in\mathbb{R}^{m}_{+}.$
Using the quadratic loss function, we derive the least-squares version of the model as follows:
$\min_{W,b,c,\xi}$ $\sum_{i=1}^{m}\|Wx_{i}+b\|_{2}^{2}+{\mathcal{C}}\sum_{i=1}^{m}\xi^{2}_{i}$ (LS-$\ell_{0}$-QSVM)
s.t. $y_{i}(\frac{1}{2}x_{i}^{\top}W{x_{i}}+b^{\top}x_{i}+c)=1-\xi_{i};\quad\forall i\in[m],$
$\|[\mbox{hvec}(W);b]\|_{0}\leq k,$
$W\in\mathbb{S}_{n},\,b\in\mathbb{R}^{n},\,c\in\mathbb{R},\,\xi\in\mathbb{R}^{m},$
To present a penalty decomposition algorithm that works for both proposed models, with different corresponding subproblems, we introduced the following unified model:
$\min_{W\in\mathbb{S}_{n},\,b\in\mathbb{R}^{n},\,c\in\mathbb{R}}$ $\sum_{i=1}^{m}\|Wx_{i}+b\|_{2}^{2}+{\mathcal{C}}\sum_{i=1}^{m}\mathrm{H}(1-y_{i}f_{W,b,c}(x_{i}))\qquad\mbox{s.t.}\qquad\|[\mbox{hvec}(W);b]\|_{0}\leq k,$
which reduces to $\ell_{0}$-QSVM for the well-received hinge loss $\mathrm{H}(t)=\max(t,0)$ and reduces to LS-$\ell_{0}$-QSVM’ for the quadratic loss $\mathrm{H}(t)=t^{2}$. Based on the notation introduced in the previous section, the above model is equivalent to:
$\min_{z\in\mathbb{R}^{\frac{n(n+1)}{2}+n},c\in\mathbb{R}}$ $\frac{1}{2}z^{\top}Gz+{\mathcal{C}}\sum_{i=1}^{m}\mathrm{H}(1-y_{i}(z^{\top}r_{i}+c))\qquad\mbox{s.t.}\qquad\|z\|_{0}\leq k.$ (1)
This unified model is still intractable, and therefore, we next apply an effective penalty decomposition method to solve it. After introducing a new variable $u$, this problem becomes:
$\min_{z,u\in\mathbb{R}^{\frac{n(n+1)}{2}+n},c\in\mathbb{R}}$ $\frac{1}{2}z^{\top}Gz+{\mathcal{C}}\sum_{i=1}^{m}\mathrm{H}(1-y_{i}(z^{\top}r_{i}+c))\qquad\mbox{s.t.}\qquad\|z\|_{0}\leq k\qquad\mbox{and}\qquad z-u=0.$
Thus, by penalizing the last constraint above, the problem 1 can be addressed by solving a sequence of penalty subproblems as follows:
$\min_{z,u\in\mathbb{R}^{\frac{n(n+1)}{2}+n},c\in\mathbb{R}}$ $q_{\rho}(z,c,u)=\frac{1}{2}z^{\top}Gz+{\mathcal{C}}\sum_{i=1}^{m}\mathrm{H}(1-y_{i}(z^{\top}r_{i}+c))+\frac{1}{2}\rho\|z-u\|_{2}^{2}\qquad\mbox{s.t.}\qquad\|u\|_{0}\leq k.$ ($P_{\rho}$)
Suppose we have a feasible point denoted as $(z,c)^{\text{feas}}$ of 1 in hand. Then, let:
$\Upsilon\geq\max\{f((z,c)^{\text{feas}}),\min_{z,c}q_{\rho^{(0)}}(z,c,u^{(0)}_{0})\}>0.$
We solve the above penalty subproblem for a fixed positive $\rho$ via a block coordinate descent method. The stopping criterion for the inner loop is the following:
$\max\left\{\frac{\|z_{l}-z_{l-1}\|_{\infty}}{\max\left(\|z_{l}\|_{\infty},1\right)},\frac{\|c_{l}-c_{l-1}\|_{\infty}}{\max\left(\|c_{l}\|_{\infty},1\right)},\frac{\|u_{l}-u_{l-1}\|_{\infty}}{\max\left(\|u_{l}\|_{\infty},1\right)},\right\}\leq\epsilon_{I}.$ (2)
And for the outer loop is:
$\|z^{(j)}-u^{(j)}\|_{\infty}\leq\epsilon_{O}.$ (3)
Therefore, we present the following penalty decomposition algorithm for solving (1).
Algorithm 1 $\ell_{0}$-Regularized QSVM Penalty Decomposition
1: Inputs: $\rho^{(0)}>0,\beta>1,k$, and $u^{(0)}_{0}$ such that $\|u^{(0)}_{0}\|_{0}\leq k$.
[…]
3:repeat
4:  $l\leftarrow 0$.
5:  repeat
6:    $(z^{(j)}_{l+1},c^{(j)}_{l+1})\leftarrow\mbox{Argmin}_{(z,c)\in\mathbb{R}^{\frac{n(n+1)}{2}+n}\times\mathbb{R}}\ q_{\rho^{(j)}}(z,c,u^{(j)}_{l})$.
7:    $u^{(j)}_{l+1}\leftarrow\mbox{Argmin}_{\{u|\|u\|_{0}\leq k\}}\ q_{\rho^{(j)}}(z^{(j)}_{l+1},c^{(j)}_{l+1},u).$
8:   $l\leftarrow l+1$.
9:  until stopping criterion (2) is met.
10:   $(z^{(j)},c^{(j)},u^{(j)})\leftarrow(z^{(j)}_{l},c^{(j)}_{l},u^{(j)}_{l})$.
11:  $\rho^{(j+1)}\leftarrow\beta\cdot\rho^{(j)}$.
12:  If $\min_{(z,c)\in\mathbb{R}^{\frac{n(n+1)}{2}+n}\times\mathbb{R}}\ q_{\rho^{(j+1)}}(z,c,u^{(j)})>\Upsilon$, then $u_{0}^{(j+1)}\leftarrow z^{\text{feas}}$. Otherwise, $u^{(j+1)}_{0}\leftarrow u^{(j)}_{l}$.
13:  $j\leftarrow j+1$.
14:until stopping criterion (3) is met.
We now turn our attention to solving the restricted subproblems within the algorithm as efficiently as possible.
Observe that the associated subproblem for $u$ is:
$\min_{u\in\mathbb{R}^{\frac{n(n+1)}{2}+n}}\ \|z-u\|_{2}^{2}\qquad\mbox{s.t.}\qquad\|u\|_{0}\leq k;$ (4)
which obtains a solution
$u=[z_{\mathcal{J}};0],$ (5)
where $\mathcal{J}$ contains indices of $k$ largest components of $z$ in absolute value, that is, Step 7 of Algorithm 1 has a closed-form solution. Next, we start discussing how to tackle the associated subproblem to $z$ for each of the loss functions.
3.1 $\ell_{0}$-Regularized Quadratic Surface SVM Model
For the hinge loss function $\mathrm{H}(t)=\max(0,t)$, the sparse quadratic model 1 reduces to $\ell_{0}$-QSVM, which after using the notation in Section 2 is equivalent to:
$\min_{z,c,\xi}$ $\frac{1}{2}{z}^{\top}Gz+{\mathcal{C}}\sum_{i=1}^{m}\xi_{i}$ ($\ell_{0}$-QSVM’)
s.t. $y_{i}(z^{\top}r_{i}+c)\geq 1-\xi_{i};\quad\forall i\in[m],$
$\|z\|_{0}\leq k.$
$z\in\mathbb{R}^{\frac{n(n+1)}{2}+n},c\in\mathbb{R},\xi\in\mathbb{R}^{m}_{+}.$
Recalling $P_{\rho}$, the subproblem associated to $z$ for the hinge loss function becomes:
$\min_{z\in\mathbb{R}^{\frac{n(n+1)}{2}+n},c\in\mathbb{R}}$ $\frac{1}{2}z^{\top}Gz+{\mathcal{C}}\sum_{i=1}^{m}\max(1-y_{i}(z^{\top}r_{i}+c),0)+\frac{1}{2}\rho\|z-u\|_{2}^{2},$
which is equivalent to the following feasible and bounded below the convex quadratic optimization problem:
$\min_{z,c,\xi}$ $\frac{1}{2}{z}^{\top}(G+\rho I)z-\rho u^{\top}z+{\mathcal{C}}\sum_{i=1}^{m}\xi_{i}$ ($P^{h}_{\rho,(z,c,\xi)}$)
s.t. $y_{i}(z^{\top}r_{i}+c)\geq 1-\xi_{i};\quad\forall i\in[m]$
$z\in\mathbb{R}^{\frac{n(n+1)}{2}+n},c\in\mathbb{R},\xi\in\mathbb{R}^{m}_{+},.$
Even though a standard solver can be used for this problem, standard practice in the SVM literature encourages applying the strong duality theorem in convex programming. Thus, let us bring the Lagrangian function of this model:
$L(z,c,\xi,\alpha)$ $=$ $\frac{1}{2}{z}^{\top}(G+\rho I)z-\rho u^{\top}z+{\mathcal{C}}\sum_{i=1}^{m}\xi_{i}-\sum_{i=1}^{m}\alpha_{i}(y_{i}(z^{\top}r_{i}+c)-1+\xi_{i})$
$=$ $\frac{1}{2}{z}^{\top}(G+\rho I)z-z^{\top}\Big{(}\sum_{i=1}^{m}\alpha_{i}y_{i}r_{i}+\rho u\Big{)}+\sum_{i=1}^{m}({\mathcal{C}}-\alpha_{i})\xi_{i}-c\sum_{i=1}^{m}\alpha_{i}y_{i}.$
Therefore, the dual problem of $P^{h}_{\rho,(z,c,\xi)}$ becomes:
$\min_{\alpha\in\mathbb{R}^{m}}$ $-\sum_{i=1}^{m}\alpha_{i}+\frac{1}{2}\Big{(}\sum_{i=1}^{m}y_{i}\alpha_{i}r_{i}+\rho u\Big{)}^{\top}(G+\rho I)^{-1}\Big{(}\sum_{j=1}^{m}y_{j}\alpha_{j}r_{j}+\rho u\Big{)}$ ($D_{\alpha}$)
s.t. $\sum_{i=1}^{m}y_{i}\alpha_{i}=0\qquad\mbox{and}\qquad 0\leq\alpha\leq\textbf{1}{\mathcal{C}}.$
Therefore, the Karush-Kuhn-Tucker (KKT) conditions are:
[…]
which implies that, once a solution to the dual $D_{\alpha}$ is provided, a solution to $P^{h}_{\rho,(z,c,\xi)}$ can be obtained from:
[…]
Corollary 3.1.
For the hinge loss function, step 6 of Algorithm 1 can be obtained by solving $D_{\alpha}$ and then applying 6.
3.2 Least-squares $\ell_{0}$-Regularized Quadratic Surface SVM Model
For the quadratic loss function $\mathrm{H}(t)=t^{2}$, the sparse quadratic model 1 reduces to the following least-squares $\ell_{0}$-regularized kernel-free quadratic surface SVM model LS-$\ell_{0}$-QSVM, which through the notations presented in Section 2 becomes:
$\min_{z,c,\xi}$ $\frac{1}{2}{z}^{\top}Gz+{\mathcal{C}}\sum_{i=1}^{m}\xi^{2}_{i}$ (LS-$\ell_{0}$-QSVM’)
s.t. $y_{i}(z^{\top}r_{i}+c)=1-\xi_{i};\quad\forall i\in[m],$
$\|z\|_{0}\leq k,$
$z\in\mathbb{R}^{\frac{n(n+1)}{2}+n},c\in\mathbb{R},\xi\in\mathbb{R}^{m},.$
The associated subproblem to $z$ in this case is equivalent to:
$\min_{z\in\mathbb{R}^{\frac{n(n+1)}{2}+n},c\in\mathbb{R}}$ $\frac{1}{2}{z}^{\top}Gz+{\mathcal{C}}\sum_{i=1}^{m}(1-y_{i}(z^{\top}r_{i}+c))^{2}+\frac{1}{2}\rho\|z-u\|^{2}_{2}.$ ($P_{(z,c)}^{q}$)
By letting $A=[r_{1},r_{2},\dots,r_{m}]^{\top}$ and $D={\rm{\mbox{\bf{diag}}}}(y_{1},y_{2},\dots,y_{m})$, the objective function of the latter is equal to $T=\frac{1}{2}z^{\top}(G+\rho I)z+{\mathcal{C}}\|\textbf{1}-D(Az+c\textbf{1})\|_{2}^{2}-\rho u^{\top}z+\rho\|u\|_{2}^{2}$. Hence,
$\frac{\partial T}{\partial z}$ $=$ $(G+\rho I)z-2{\mathcal{C}}A^{\top}D(\textbf{1}-D(Az+c\textbf{1}))-\rho u$
$\frac{\partial T}{\partial c}$ $=$ $-2{\mathcal{C}}\textbf{1}^{\top}D(\textbf{1}-D(Az+c\textbf{1}))$
which can be integrated into the following:
$\begin{bmatrix}G+\rho I+2{\mathcal{C}}A^{\top}D^{2}A&2{\mathcal{C}}A^{\top}D\textbf{1}\\ 2{\mathcal{C}}\textbf{1}^{\top}D^{2}A&2{\mathcal{C}}\textbf{1}^{\top}D^{2}\textbf{1}\end{bmatrix}\begin{bmatrix}z\\ c\end{bmatrix}=\begin{bmatrix}2{\mathcal{C}}A^{\top}D\textbf{1}+\rho u\\ 2{\mathcal{C}}\textbf{1}^{\top}D\textbf{1},\end{bmatrix}$
which has a positive definite coefficient matrix such that:
$\begin{bmatrix}z\\ c\end{bmatrix}=\begin{bmatrix}G+\rho I+2{\mathcal{C}}A^{\top}D^{2}A&2{\mathcal{C}}A^{\top}D\textbf{1}\\ 2{\mathcal{C}}\textbf{1}^{\top}D^{2}A&2{\mathcal{C}}\textbf{1}^{\top}D^{2}\textbf{1}\end{bmatrix}^{-1}\begin{bmatrix}2{\mathcal{C}}A^{\top}D\textbf{1}+\rho u\\ 2{\mathcal{C}}\textbf{1}^{\top}D\textbf{1},\end{bmatrix}$ (7)
Corollary 3.2.
For the quadratic loss function, step 6 of Algorithm 1 is obtained from (7).
To conclude, we present two sparse $\ell_{0}$-norm, kernel-free quadratic surface SVMs that employ hinge and quadratic loss functions, respectively. Because the presence of a sparsity constraint renders these formulations NP-hard, an efficient solution method is essential. We place both models within the unified framework (1) and develop a penalty decomposition algorithm that introduces an auxiliary variable $u$. The associated subproblem (4) admits a simple closed-form solution (5). For the hinge-loss-based model ($\ell_{0}$-QSVM), the $z$-subproblem is solved most effectively via convex duality, with the primal variables recovered through (6). In the quadratic-loss-based model (LS-$\ell_{0}$-QSVM), the $z$-subproblem itself has the closed-form solution given in (7). Consequently, each iteration of our algorithm remains computationally efficient while enforcing exact sparsity.
3.3 Convergence of $\ell_{0}$-Regularized QSVM Penalty Decomposition Algorithm
While convergence guarantees are well established for convex models, the presence of a nonconvex and nonsmooth $\ell_{0}$-norm constraint introduces significant theoretical and computational challenges. The proposed models belong to the class of combinatorial optimization problems, for which finding global minimizers is generally intractable. To address this, we adopt a penalty decomposition framework whose subproblems are either convex or admit closed-form solutions, depending on the choice of loss function. For the hinge loss, the main subproblem reduces to a linearly constrained quadratic program, which can be efficiently solved via duality theory. In the case of quadratic loss, the main subproblem simplifies to solving a system of linear equations, for which a closed-form solution is obtained. Specifically, Algorithm 1 is a direct application of the general penalty decomposition method proposed in [9], applied to the context of $\ell_{0}$-regularized kernel-free quadratic SVMs. This framework circumvents the combinatorial difficulty of direct $\ell_{0}$-minimization by solving a sequence of tractable subproblems, each enforcing sparsity in a relaxed yet principled manner.
In the subsequent analysis, we demonstrate that Algorithm 1 converges to a solution satisfying the Lu-Zhang stationarity conditions–a generalized notion of first-order optimality tailored to nonconvex problems with cardinality constraints. This concept extends the classical KKT framework and is particularly suitable for settings where conventional convex stationarity conditions fail to capture meaningful structure. To formally analyze convergence, we consider a general class of sparse optimization problems with structural constraints, as described in problem (P{P}). These problems are motivated by applications in signal recovery, image reconstruction, and data compression [30] and can be formulated as:
$\min_{{v}\in\mathbb{R}^{\hat{\jmath}}}\,{\varphi}({v})\qquad\text{subject to}\qquad{\psi}({v})\leq 0,\quad{\chi}({v})=0,\quad\|{v}\|_{0}\leq k,$ (P{P})
where $\varphi$}, $\psi$}, and $\chi$} are continuously differentiable functions, though not necessarily convex, and $k<\hat{\jmath}$ denotes the desired sparsity level.
A point v{v} satisfying ${\psi}({v})\leq 0$ and ${\chi}({v})=0$ is called a Lu-Zhang stationary point for problem (P{P}) if there exists an index set $\mathcal{L}\subseteq[\hat{\jmath}$] with $|\mathcal{L}|=k$ such that ${v}_{j}=0$ for all $j\in\mathcal{L}^{c}:=[\hat{\jmath}$$\setminus\mathcal{L}$, and there exist multipliers ${\lambda}\in\mathbb{R}^{\tilde{q}}$ and ${\mu}\in\mathbb{R}^{\check{r}}$ satisfying:
$\begin{cases}\nabla{\varphi}({v})-{\lambda}^{\top}\nabla{\psi}({v})-{\mu}^{\top}\nabla{\chi}({v})-{\omega}=0,\\ {\lambda}_{i}\geq 0,\quad{\lambda}_{i}\cdot{\psi}_{i}({v})=0,\quad\forall i,\quad\text{and}\quad{\omega}_{j}=0,\quad\forall j\in\mathcal{L}^{c}.\end{cases}$ (8)
While alternative stationarity concepts exist (e.g., basic feasibility [31]), Lu-Zhang stationarity provides a suitable theoretical foundation for convergence analysis in nonconvex sparse settings. It is known that basic feasible points are also Lu-Zhang stationary points, though the converse does not always hold [32].
[…]
Next, observe that Algorithm 1 is a direct application of the general penalty decomposition framework developed in [9], and thus inherits its convergence guarantees without further modification. In particular, Theorem 4.3 in [9] guarantees convergence to a Lu-Zhang stationary point, provided that Robinson’s constraint qualification holds—an assumption automatically satisfied in our models due to the linearity of their constraint sets $\mathbb{K}$. Furthermore, since the objective functions in both formulations are convex, any Lu-Zhang stationary point with full support of size $k$ is a local minimizer. Hence, Algorithm 1 produces feasible $k$-sparse solutions where the gradient of the Lagrangian vanishes on the support, certifying convergence to a Lu-Zhang stationary point of problem (1). More specifically, in solving the hinge loss model $\ell_{0}$-QSVM’, Algorithm 1 obtains a point $(z,c,\xi)\in\mathbb{R}^{\frac{n(n+1)}{2}+n}\times\mathbb{R}\times\mathbb{R}^{m}$ such that $z$ is supported on a set $\mathcal{L}\subseteq\left[\frac{n(n+1)}{2}+n\right]$ with $|\mathcal{L}|=k$, and the following conditions hold:
$\left\{\begin{array}[]{ll}Gz-\sum_{i=1}^{m}\lambda_{i}y_{i}r_{i}-{\omega}=0,&\text{where }{\omega}_{j}=0\text{ for }j\in\mathcal{L}\\[4.0pt] \sum_{i=1}^{m}\lambda_{i}y_{i}=0,&\\[4.0pt] {\mathcal{C}}-\lambda_{i}-\bar{\lambda}_{i}=0,&\forall i\in[m]\\[4.0pt] \lambda_{i}\geq 0,\quad\lambda_{i}\cdot\left(y_{i}(z^{\top}r_{i}+c)+\xi_{i}-1\right)=0,&\forall i\in[m]\\[4.0pt] \bar{\lambda}_{i}\geq 0,\quad\bar{\lambda}_{i}\cdot\xi_{i}=0,&\forall i\in[m]\\[4.0pt] y_{i}(z^{\top}r_{i}+c)\geq 1-\xi_{i},\quad\xi_{i}\geq 0,&\forall i\in[m]\end{array}\right$.
Furthermore, when solving the quadratic loss model LS-$\ell_{0}$-QSVM’, Algorithm 1 returns a point $(z,c,\xi)$ satisfying $z_{j}=0$ for all $j\in\mathcal{L}^{c}:=\left[\frac{n(n+1)}{2}+n\right]\setminus\mathcal{L}$ with $|\mathcal{L}|=k$, and the following Lu-Zhang stationarity conditions hold:
$\left\{\begin{array}[]{ll}Gz-\sum_{i=1}^{m}\mu_{i}y_{i}r_{i}-{\omega}=0,&\text{where }{\omega}_{j}=0\text{ for }j\in\mathcal{L}\\[4.0pt] \sum_{i=1}^{m}\mu_{i}y_{i}=0,&\\[4.0pt] 2{\mathcal{C}}\xi_{i}+\mu_{i}=0,&\forall i\in[m]\\[4.0pt] y_{i}(z^{\top}r_{i}+c)=1-\xi_{i},&\forall i\in[m].\end{array}\right$.
4 Numerical Experiments
In this section, we perform numerical experiments on real datasets to evaluate the performance of our proposed models $\ell_{0}$-QSVM and LS-$\ell_{0}$-QSVM when solved by Algorithm 1. We will also compare their accuracy and ability to recover true underlying sparsity patterns against the benchmark methods QSVM and $\ell_{1}$-QSVM.
We implemented all QSVM, $\ell_{1}$-QSVM, $\ell_{0}$-QSVM and LS-$\ell_{0}$-QSVM models in Python. We evaluated the accuracy of these models on several balanced datasets from the UCI repository, including Iris, Glass Identification, Raisin, Wine Quality, Hepatitis C Virus for Egyptian patients (HCV), Yeast, Wholesale Customers, Heart, Ionosphere, Sonar, Brain, and Bill. These datasets are widely used benchmarks for evaluating the performance of newly developed models in the literature. They encompass both binary and multi-class label variables. For multi-class datasets, we utilized the one-vs-rest strategy combined with majority voting to determine the predicted class.
We performed hyperparameter tuning using random search. For the kernel-free models $\ell_{1}$-QSVM, $\ell_{0}$-QSVM, and LS-$\ell_{0}$-QSVM, we tuned the regularization parameter $\mathcal{C}$} over a logarithmic scale from $10^{-2}$ to $10^{2}$. In addition, for the sparse models $\ell_{0}$-QSVM and LS-$\ell_{0}$-QSVM, we tuned the sparsity level $k$ over the range $1$ to $2n$, where $n$ denotes the number of original features. We evaluated performance under both hinge loss and quadratic loss formulations.
We emphasize that while a generic quadratic model typically requires $\frac{n(n+1)}{2}+n+1$ parameters, for the implemented experiments, our sparse models are set to use at most $2n+1$, resulting in a significantly more compact representation. Despite this reduction in model complexity, experimental results demonstrate that our approach achieves competitive predictive performance, even in a few cases where it does not outperform existing methods. These findings underscore overfitting as a key concern in kernel-free quadratic models, one that necessitates careful mitigation through principled sparsity-inducing strategies.
To evaluate model performance, we employed $5$-fold cross-validation. Each dataset was partitioned into five folds, with each fold serving as the test set in turn, while the remaining four folds were further split into three for training and one for validation. Testing was conducted using models tuned with the best hyperparameters selected from the validation set. For each fold, we sampled 100 different hyperparameter configurations for tuning. This process produced five accuracy values for each dataset, from which we calculated the mean and standard deviation, as reported in Table 1.
Table 1: Test Accuracy Comparisons
Dataset QSVM $\ell_{1}$-QSVM $\ell_{0}$-QSVM LS-$\ell_{0}$-QSVM
Iris 0.953 $\pm$ 0.061 0.953 $\pm$ 0.077 0.953 $\pm$ 0.034 0.960 $\pm$ 0.049
Glass 0.631 $\pm$ 0.041 0.640 $\pm$ 0.033 0.650 $\pm$ 0.051 0.656 $\pm$ 0.062
Raisin 0.862 $\pm$ 0.030 0.863 $\pm$ 0.029 0.859 $\pm$ 0.030 0.851 $\pm$ 0.031
Wine 0.955 $\pm$ 0.022 0.961 $\pm$ 0.029 0.977 $\pm$ 0.021 0.961 $\pm$ 0.034
HCV 0.936 $\pm$ 0.013 0.939 $\pm$ 0.013 0.941 $\pm$ 0.005 0.924 $\pm$ 0.019
Yeast 0.610 $\pm$ 0.024 0.604 $\pm$ 0.030 0.590 $\pm$ 0.027 0.567 $\pm$ 0.028
Wholesale 0.902 $\pm$ 0.023 0.905 $\pm$ 0.027 0.900 $\pm$ 0.022 0.893 $\pm$ 0.016
Heart 0.816 $\pm$ 0.034 0.796 $\pm$ 0.102 0.822 $\pm$ 0.022 0.832 $\pm$ 0.034
Ionosphere 0.934 $\pm$ 0.028 0.934 $\pm$ 0.032 0.866 $\pm$ 0.023 0.894 $\pm$ 0.025
Sonar 0.836 $\pm$ 0.065 0.789 $\pm$ 0.043 0.774 $\pm$ 0.016 0.764 $\pm$ 0.016
Brain 0.938 $\pm$ 0.032 0.947 $\pm$ 0.011 0.935 $\pm$ 0.034 0.947 $\pm$ 0.018
Bill 0.960 $\pm$ 0.083 0.967 $\pm$ 0.083 0.972 $\pm$ 0.071 0.959 $\pm$ 0.056
Next, to illustrate the strength of the $\ell_{0}$ norm in identifying effective sparse patterns that perform well in practice, we visualize these patterns in both 2D and 3D. Figure 1 demonstrates that $\ell_{0}$-QSVM retains high performance even in lower-dimensional spaces that are visualizable. The figure shows the decision boundaries of $\ell_{0}$-QSVM for the Iris dataset with $k=2$. In Figure 2, the Wine Quality dataset, the decision boundary appears as a hyperplane separating the classes, whereas in the Raisin dataset, it is depicted as a curved 3D manifold enclosing one class and excluding another. These examples highlight the potential of sparse $\ell_{0}$ QSVMs for data visualization, enabling projection into two- or three-dimensional feature spaces.
Figures 3 and 4 illustrate a comparison of the weight matrices $W$ and vectors $b$ for $\ell_{1}$-QSVM and $\ell_{0}$-QSVM when applied to the Iris and Glass datasets. The weight matrices were obtained using the optimal hyperparameters for each model, which were determined through validation on a randomly selected 20% subset of the Iris and Glass datasets. As evident from the figures, the $W$ matrices and $b$ vectors for $\ell_{0}$-QSVM exhibit significantly greater sparsity compared to those of $\ell_{1}$-QSVM. This sparsity substantially enhances the interpretability of our classifiers, surpassing what is suggested by their $\ell_{1}$-norm counterparts. Nevertheless, despite this increased sparsity, Table 1 shows that the accuracy of $\ell_{0}$-QSVM is either higher than or comparable to that of QSVM.
Figure 1: $\ell_{0}$-QSVM Classifier for $k=2$ on Iris dataset
Figure 2: Surfaces obtained by $\ell_{0}$-QSVM with $k=3$ on Raisin and Wine datasets
Figure 3: Magnitudes of optimal parameters (weight matrix $W$ and vector $b$) of classifiers found by $\ell_{1}$-QSVM (a) and $\ell_{0}$-QSVM (b) for the Iris dataset.
Figure 4: Magnitudes of optimal parameters (weight matrix $W$ and vector $b$) of classifiers found by $\ell_{1}$-QSVM (a) and $\ell_{0}$-QSVM (b) for the Glass dataset.
Figure 5 illustrates the accuracy of $\ell_{0}$-QSVM across different values of $k$. As observed, these models often achieve their full predictive power with relatively small values of $k$, while maintaining the flexibility of models incorporating second-order features. Unlike traditional kernel methods that transform the data into a higher-dimensional feature space, $\ell_{0}$-QSVM operates directly in the original feature space, enhancing interpretability. This quality is particularly valuable in domains like healthcare, where model interpretability is a critical requirement.
We emphasize that the strength of our model lies not necessarily in outperforming every competitor on all datasets, but in providing an explainable model that achieves performance close to that of well-established competitors commonly used in practice. As Figure 5 reveals, in our approach, the maximum reachable accuracy is usually acquired with a small value of $k$, making the resulting decision boundary more interpretable and easier to understand. As shown in Figure 5, our approach typically achieves the maximum attainable accuracy with a small value of $k$. This enhances the interpretability of the resulting decision boundary.
Figure 5: Accuracy vs sparsity level $k$
5 Conclusion
This paper introduces $\ell_{0}$-regularized kernel-free quadratic surface SVM models to address challenges such as overfitting, interpretability, and the complexity of kernel-free quadratic surface machines in binary classification tasks. By incorporating a sparsity constraint on model parameters, these methods effectively reduce over-parameterization while enhancing generalization and sparsity. However, these models are inherently intractable and NP-hard, underscoring the need for an effective algorithm to solve them. To this end, we developed a penalty decomposition algorithm that leverages duality theory and provides closed-form solutions for key subproblems. Numerical experiments conducted on real-world datasets demonstrate that the proposed models outperform standard linear and quadratic SVM models and $\ell_{1}$-regularized counterparts in detecting true underlying sparse patterns while maintaining high classification accuracy. These results underscore the practical advantages of the $\ell_{0}$-regularized approach in applications requiring interpretable and sparse decision boundaries.
The proposed framework holds significant promise for fields where transparency and interpretability are critical, such as healthcare and finance. In addition to its standard formulation, the twin version of the proposed models enhances efficiency by solving smaller, decoupled subproblems, making it suitable for high-dimensional data. Moreover, the framework’s ability to handle imbalanced scenarios is a notable strength, as it effectively incorporates sparsity constraints to mitigate the dominance of majority classes, ensuring better representation of minority classes. Future research could explore scalability to larger datasets, adaptive methods for hyperparameter tuning, extensions to multi-class classification problems, strategies to improve twin formulations, and techniques to address imbalance in more complex real-world datasets, thereby further enhancing the applicability and robustness of the models across diverse settings.
Declaration of Generative AI and AI-assisted Technologies in the Writing Process
During the preparation of this work, the authors used ChatGPT to improve language and readability with extreme caution. After using this tool, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication.